OpenStudy (anonymous):

Which of the following equations is equivalent to y = lnx ? a. x = e^y b. y = e^x c. y = x^e d. x = y^e

OpenStudy (bibby):

Remember the property of logs. (natural logs anyways) a = e^b if you take the ln of both sides of that, you end up with ln(a) = b*ln(e)

OpenStudy (anonymous):

huh?

OpenStudy (bibby):

whoops misread the question. what is ln(e^x)?

OpenStudy (anonymous):

I dont understand

OpenStudy (bibby):

Then go learn the properties of natural log and come back.

OpenStudy (anonymous):

Take the ln of each of the answers and one will be what you seek.

OpenStudy (bibby):

the third property in this image in particular

OpenStudy (anonymous):

is it d?

OpenStudy (solomonzelman):

the first question you asked. y = lnx ? a. x = e^y b. y = e^x c. y = x^e d. x = y^e \[y=lnx~~~~~~~~~->~~~~~~y=\log_ex.\]\[formula:~~~~~~~~~\log_ab=c~~~~~~->~~~~~~a^c=b\]

OpenStudy (anonymous):

so you're saying it's y = e^x ?

OpenStudy (bibby):

if you take the ln of both sides of that you get ln y = x ln e and ln e simplifies to 1.

OpenStudy (anonymous):

where did the e come from?

OpenStudy (anonymous):

try taking the ln of both sides of a)

OpenStudy (bibby):

y = e^x ? ln(y) = ln(e^x)

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

\[\ln~(anything)=\log_e(anything)\]

OpenStudy (bibby):

she's clearly interested in learning about the math behind this

OpenStudy (solomonzelman):

yes, I see this.

OpenStudy (solomonzelman):

y = lnx is just saying the same thing as y = log(base e) x and you were right, it is e^y=x